OpenStudy (anonymous):

Rewrite the quadratic function f(x) = 2x^2+8x-3 in vertex form a(x-h)^2+k

OpenStudy (anonymous):

I can see that I would factor out a 2 to start it and so I would have 2(x^2+4x) - 3 but I don't know what to do after that

OpenStudy (anonymous):

f(x) = 2(x^2+4/2x)-3 b = 4/2 * 2 = 8/2 = 4 b^2 = 16 Is this on the right track?

OpenStudy (anonymous):

@iambatman @Kainui

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Do you know how to complete the square? f(x) = 2x^2+8x-3 f(x) = 2(x^2+4x) - 3 f(x) = 2(x^2+4x +4) - 3 - 2(4) Then finish

OpenStudy (anonymous):

I just watched a video on how to do it so I think this may be right

OpenStudy (anonymous):

f(x) = 2(x^2 + 4/2x + 16 - 16) - 3 f(x) = 2(x+4)^2 - 32 - 3 f(x) = 2(x+4)^2 - 35

OpenStudy (anonymous):

but when I put that in wolfram alpha, I get 2x^2+16x-3 instead of 2x^2+8x-3

OpenStudy (anonymous):

hmm, I got 4/2x + 16 in one of my steps

OpenStudy (anonymous):

Would it not be 4^2 = 16 and not +4?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Divide the middle coefficient (here it's 4) by 2, and square it: f(x) = 2x^2+8x-3 f(x) = 2(x^2+4x) - 3 f(x) = 2(x^2+4x +4) - 3 - 2(4) f(x) = 2(x+2)^2 - 3 - 2(4)

OpenStudy (anonymous):

f(x) = 2(x+2)^2 - 11

OpenStudy (anonymous):

How did you get f(x) = 2(x+2)^2?

OpenStudy (agent0smith):

By factoring. x^2 + bx + (b/2)^2 factors as (x + b/2)^2